Andretti Company has a single product called a Dak. The company normally produces and sells 89,000 Daks each year at a sellin4. Due to a strike in its suppliers plant, Andretti Company is unable to purchase more material for the production of Daks.
Answer #1
Contribution margin
selling price per unit 60
less Variable expenses
direct materials 6.5
direct labor 9
Variable manufacturing overhead 3.1
variable selling expense 3.7 22.3
Contribution margin per unit 37.7
Req 1A increased sales in units (89000*20%)= 17800
contribution margin per unit 37.7
incremental contribution margin 671060
less added fixed selling expense 140,000
incremental net operarting income 531,060
1-b) Yes
Req 2 Break even price per unit
Variable manufacturing cost per unit 18.6
Shipping cost 2.5
import duties 4.7
permits &licences 0.7
Break even price per unit 26.5 answer
Req 3 Relevant unit cost $3.70 per unit
4) Foregone contribution margin (3708*37.7) 139791.60
total avoidable fixed cost
fixed manufacturing overhead cost (445000*2/12)*70% 51916.67
fixed selling cost (267000*2/12)*20% 8900 60816.67
Financial disadvantage -78974.93
89000*2/12*25%= 3708.333 units
No
5) Variable manfuacturing costs 18.6
fixed manufacturing overhead cost (5*30%)= 1.5
variable selling expense 3.7*1/3 1.23
total costs avoided 21.33